Using Excel: You were asked by the Manager of Engineering to propose a solution for a current production line problem. After analyzing the issues you came up with two solutions, both of which will solve the problem for the next five years. Solution A would initially cost $90,000, have annual O&M costs of $22,000 and will generate annual savings of $48,000, while solution B will need an initial $62,000, $17,000 for annual O&M costs, and will generate annual savings of $36,000. Both will have salvage values, $15,000 for solution A, and $10,000 for B. Your company's marginal income tax rate is 40%, and its MARR is 10%. The proposed equipment for both options is subject to a five-year MACRS property class.

a.What value of MARR would make each solution break even?

b.What would be the amount of additional revenue that Solution A should generate to make the Manager of Engineering indifferent to the choice between the two options?
Answer #1

Answer :

(a).

Calculation of the amount of MAAR at which solution Break Even
Sl No. Particulars Solution A Solution A
1. Initial cost $ 90,000 $ 62,000
2. Annual O$M Costs $ 22,000 $ 17,000
3. Annual Savings $ 48,000 $ 36,000
4. Salvage value $ 15,000 $ 10,000
5. Life in years 5 5
6. Depreciation per annum 90000/5 = 18000 62000/5 =12400
7. Net savings 48000-22000 = 26000 36000-17000 =19000
8. Net Savings less Depreciation 26000-18000 =8000 19000-12400 = 6600
9. Less Tax @40 8000*0.4 = 22800 6600*0.4 = 2640
10. Net savings after tax but before depreciation 4800+18000 = 22800 3960+12400 = 16360
11. Annuity Discounting factor MARR @ 10% for 5 years 3.791 3.791
12. 5th Year present value dis factor @ 10% 0.621 0.621
13. Net Savings after tax but before depreciation @10% 3.791*22800 = 86435 3.791*16360 = 62021
14. salvage value of 5th year @ 10% 0.621*15000 = 9315 0.621*15000 = 6210
15. Total Savings 86435+9315 =95750 62021+6210=68231
16. Net Present value 95750-90000 5750 68231-62000=6231
17. NPV as a % od initial cost 5750/90000*100 = 6.39% 6231/62000*100=10.05%
18. Internal rate of return- Break even 10%+6.39%=16.39% 10%+10.05%=20.05%

NOTES

The standard method of depreciation for federal income tax purposes is called the modified Accelerated cost recovery System or MACRS. Essentially a MACRS depreciation schedule will begin with a declining balance method then switch to a straight line schedule to finish the schedule.

(b).

Project A Yeari Year5 YearO Year2 Year3 Year4 particulars Cost Annual cost Saving Net saving(Saving-cost)(A)-90000 Salvage TaProject B Yearl Year2 Year4 YearO particulars Cost Annual cost Saving Net saving(Saving-cost) 62,000.00 19,000.00 19,000.00 1

Additional Revenue
Year 0 Year 1 Year 2 Year 3 Year 4 Year 5
Revenue from A 26,000.00 26,000.00 26,000.00 26,000.00 26,000.00
Revenue from B 19,000.00 19,000.00 19,000.00 19,000.00 19,000.00
Revenue income of A over B 7,000.00 7,000.00 7,000.00 7,000.00 7,000.00
Revenue from A is Already over and above B